Ahydraulic press has pistons of area 0.006 m² and 0.35 m², respectively. A force of 25 N is applied to the smaller one. Find the pressure exerted on the smaller piston and on the big piston. What is the lifting force on the bigger piston?


1
Expert's answer
2022-03-30T14:09:47-0400
"\frac{F_1}{F_2}=\frac{A_1}{A_2}"

"F_2=F_1\frac{A_2}{A_1}=25*\frac{0.35}{0.006}=1458\:\rm N"